OpenStudy (anonymous):

What is the equation of the line, in slope-intercept form, that passes through (3, -1) and (-1, 5)?

OpenStudy (anonymous):

m = (y2 - y1)/(x2 - x1) = (-1 - 5)/(3 - -1) = -6/(3 + 1) = -6/4 = -3/2 The slope is -3/2, so the equation so far is y = (-3/2)x + b. Now solve for b by substituting one of the points for x and y into the equation. I'll use point (3, -1): y = (-3/2)x + b -1 = (-3/2)(3) + b -1 = -9/2 + b -2/2 = -9/2 + b 7/2 = b ANSWER: y = (-3/2)x + 7/2